OpenStudy (anonymous):

Evaluate the integral of the vector field F=(x/yz)i + (y/xz)j + (z/xy)k along the line r = (cost t) i + (sin t) j + (cos t) k where t varies from pi/6 to pi/3.

OpenStudy (anonymous):

what have you tried ?

OpenStudy (experimentx):

Find dr first \[ \int_{\pi \over 6}^{\pi \over 3} \vec F \cdot d\vec r \] rest should be easy

OpenStudy (anonymous):

F • dr = F(r(t)) • r'(t) dt F(r(t)) = csc(t), sec(t)tan(t), csc(t) r'(t) = -sin(t), cos(t), -sin(t)

OpenStudy (anonymous):

yep.. so F dot dr/dt = tan(t) -2

OpenStudy (anonymous):

integrate and evaluate

OpenStudy (anonymous):

Integrate tan(t)-2 and then evaluate for the values of t given?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it -2 t-log(cos(t))

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Great! So where do I input the two t values in this equation?

OpenStudy (anonymous):

they're the limits of integration...

OpenStudy (anonymous):

\[\huge \int\limits_{\frac{ \pi }{6 }}^{\frac{ \pi }{ 3 }} (\tan t -2 )dt\]

OpenStudy (anonymous):

So im calculating [-2(pi/3)-log(Cos(pi/3))]-[-2(pi/6)-log(Cos(pi/6))]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can I calculate this straight out on a calculator or should I show working?

OpenStudy (anonymous):

go with the exact answer...

OpenStudy (anonymous):

Just checking to see what is good practice. I get -0.808 that doesnt look right does it?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

I know what I did I type in log on calculator when i should have used ln. Rookie mistake! Answer is -0.497891

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

Cheers Algebraic you've been great!!

OpenStudy (anonymous):

1/2*ln(3) - pi/3

OpenStudy (anonymous):

Whats that?

OpenStudy (anonymous):

the exact answer

OpenStudy (anonymous):

Oh I shouldnt solve completely to a numeric value?

OpenStudy (anonymous):

not usually.

OpenStudy (anonymous):

Ok thanks for that.